OpenStudy (maleena_1):

A triangle has side lengths of 18 cm, 80 cm and 81cm, Classify it as acute, obtuse, or right

OpenStudy (mathstudent55):

Do you know the Pythagorean theorem? If it's a right triangle, it must make the Pythagorean formula true. \(a^2 + b^2 = c^2\) Try it.

OpenStudy (daniel.ohearn1):

First I'd recommend drawing in relatively close to scale

OpenStudy (maleena_1):

it didn't work

OpenStudy (mathstudent55):

\(18^2 + 80^2 = 81^2\) \(324 + 6400 = 6724 = 82^2\) \(18^2 + 80^2 = 82^2\)

OpenStudy (mathstudent55):

With sides 18 and 80, 81 does not make a right triangle, but a third side of 82 would.

OpenStudy (maleena_1):

so its acute?

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

Correct.

OpenStudy (maleena_1):

ok thanks I think I got it now